OpenStudy (anonymous):

Arlene made a necklace using red and silver beads in the ratio 4 to 1. The necklace used sixty beads altogether. How many silver beads did Arlene use to make the necklace? twelve fifteen thirty-six forty-eight

OpenStudy (anonymous):

12.. cuz she used 60 total and the ratio is 4:1 red an silver respectively so she used 5 beads 12 times (5*12=60) and 1 silver bead each time (12*1=12) and 4 red beads each time (4*12=48) so 12 silver and 48 reds = 60 beads total